The lever arm hypothesis for muscle arm contraction posits that actin sliding velocity should be proportional to the myosin lever arm length. A wild-type myosin lever arm is 80 Å and generates an actin sliding velocity of 2.5 μm/s. You generate a myosin mutant with 42 amino acids deleted, resulting in a lever arm of 60 Å. You measure the resulting actin sliding velocity to be 1.875 μm/s. Based on the lever arm hypothesis, what actin sliding velocity should you expect for a longer myosin mutant with a lever arm of 100 Å.

A) 2.875 μm/s

B) 3 μm/s

C) 3.125 μm/s

D) 4.25 μm/s

E) 5 μm/s
Answer #1

Answer- the correct option is (C) that is 3.125um/s because

80A of myosin needs velocity of = 2.5 um/s

1 A of myosin needs velocity of =2.5/80

100A of myosin needs velocity of = 2.5×100/80

= 2.5×100/80 = 3.125

So the answer is 3.125um/s